OpenStudy (anonymous):

T(h)=√h/2 where H is the maximum height If a person jumps 5 feet off the ground, determine how many seconds the person is in the air

OpenStudy (anonymous):

\[T(5)=\sqrt{5}/2 s\]

OpenStudy (anonymous):

is that how I substitute the height h into the function and evaluate the expression

OpenStudy (anonymous):

the person jumps 5 feets, thats the hight, so it is h, H is max hight but it is not used in your equation, maths is case sensitive, H != h, the equationf or hight u wrote is not correct, suppose it was, then thats how u use

OpenStudy (anonymous):

so if the total hang time T in seconds is given by T(h)=√h/2

OpenStudy (anonymous):

T(5)=\sqrt{5}/2 s

OpenStudy (anonymous):

i got 1.12 seconds

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

@onaogh but you did not write it like that